Find number of students securing marks >104, between 155-165, <55 from the following class data.




Class(marks) frequency (students)




0-20 210




20-40 115




40-60 130




60-80 220




80-100 120




100-120 101




120-140 120




140-160 144




160-180 132




180-200 190

1
Expert's answer
2022-02-16T11:34:23-0500

Marks frequency(f) cumulative frequency(cf)

0-20 210 210

20-40 115    325

40-60 130 455

60-80 220 675

80-100 120 795

100-120  101 896

120-140 120  1016

140-160 144 1160

160-180 132 1292

180-200   190 1482

We are required to find the number of students securing marks greater than 104. To do so, we need to determine the "i^{th}" percentile such that "P_i=104". First is to determine the value of "i" which in turn will be used to find the position of the "104^{th}" mark. The position of the "104^{th}" mark is its cumulative frequency. We proceed as follows,

"P_i={l+({i\times n\over100}-cf)\times{c\over f}}" where, "n=1482"

"l" is the lower class boundary of the class with 104 marks

"cf" is the cumulative frequency of the class preceding the class consisting of the mark, 104

"c" is the width of the class with the 104th mark

"f" is the frequency of the class with the 104th mark

Therefore,

"P_i=100+(14.82i-795)\times {20\over101}=104"

So,

"20.2=14.82i-795\implies 14.82i=815.2\implies i=55.0067476"

The mark, 104 is approximately the "55^{th}" percentile. Its position is, "{i\times n\over 100}={55\times1482\over 100}=815.2\approx 816"

Now, the cumulative frequency of the "104^{th}" mark is approximately 816.

The number of students securing marks greater than 104 is "1482-816=666"

Therefore, the number of students securing marks greater than 104 is 666.



"b)"

To find the number of students securing marks between 155 and 165, we find the cumulative frequencies for both scores and then determine the difference of their frequencies. That is, "cf_{165}-cf_{155}".


The cumulative frequency of the "155^{th}" mark.

We determine the value "i" such that, "P_i=155" where "P_i" is the "i^{th}" percentile given as,

"Pi=l+({i×n\over100}−cf)\times {c\over f}" where, "n=1482"

"l" is the lower class boundary of the class with the "155^{th}" mark.

"cf" is the cumulative frequency of the class preceding the class with the "155^{th}" mark.

"c" is the width of the class with the 155th mark

"f" is the frequency of the class with the "155^{th}" mark

Now,

"P_i=140+(14.82i-1016)\times {20\over 144}=155"

So,

"108=14.82i-1016\implies i=75.8434548\approx76"

The score of 155 is the "76^{th}" percentile. Its cumulative frequency is, "{76\times 1482\over 100}=1124"

Therefore, the number of students securing below 155 marks is 1124


 The cumulative frequency of the "165^{th}" mark.

We determine the value "i" such that, "P_i=165" where "P_i" is the "i^{th}" percentile given as,

"Pi=l+({i×n\over100}−cf)\times {c\over f}" where, "n=1482"

"l" is the lower class boundary of the class with the "165^{th}" mark.

"cf" is the cumulative frequency of the class preceding the class with the "165^{th}" mark.

"c" is the width of the class with the "165^{th}" mark

"f" is the frequency of the class with the "165^{th}" mark

Now,

"P_i=160+(14.82i-1160)\times {20\over 132}=165"

So,

"33=14.82i-1160\implies i=80.4993252\approx81"

The score of 165 is the "81^{st}" percentile. Its cumulative frequency is, "{81\times 1482\over 100}=1193"

Therefore, the number of students securing below 165 marks is 1193


Therefore, "cf_{165}=1193" and "cf_{155}=1124" . The number of students securing marks between 155 and 165 is 1193-1124=69 students.


"c)"

The cumulative frequency of the "55^{th}" mark.

We determine the value "i" such that, "P_i=55" where "P_i" is the "i^{th}" percentile given as,

"Pi=l+({i×n\over100}−cf)\times {c\over f}" where, "n=1482"

"l" is the lower class boundary of the class with the "55^{th}" mark.

"cf" is the cumulative frequency of the class preceding the class with the "55^{th}" mark.

"c" is the width of the class with the "55^{th}" mark

"f" is the frequency of the class with the "55^{th}" mark

Now,

"P_i=40+(14.82i-325)\times {20\over 130}=55"

So,

"97.5=14.82i-325\implies i=28.5087719\approx29"

The score of 55 is the "29^{th}" percentile. Its cumulative frequency is, "{29\times 1482\over 100}=422.5\approx 423"

Therefore, the number of students securing below 55 marks is 423.